# BlockFUL: Enabling Unlearning in Blockchained Federated Learning

### v2

@@ 9                   Re-training according to (18);
11            Send model_update message($w^{\prime}_{i}$); // broadcast to the network.
12             $v^{\prime}_{i,j}$leftarrow Committee.Overwrite($\mathcal{T}(i,j)$, $\mathcal{T}^{\prime}(i,j)$, $v_{i,j},CH_{sk}$);
13             $R^{\prime}_{curr}$leftarrow Committee.Overwrite($version$,
14            $new$ $version$, $R_{curr},CH_{sk}$);
@@
17                   if $W^{\prime},v^{\prime}_{i,j},R^{\prime}_{curr}$ $is$ $invalid$ then
@@ 20                         Update the corresponding parameters in IPFS;
26            Clients($N^{1}\left(W\right)$) $\leftarrow$ model_update message($W^{\prime}$); // Parallel transmission.
@@
Algorithm 2 Sequential unlearning (re-training)
V-C Cost Analysis of Blockchain Operations
@@
This subsection analyzes the computational overhead of parallel unlearning and sequential unlearning paradigms, as well as block updating on the live chain. The update is executed every time a node is traversed, regardless of whether or not unlearning has been performed on the node. The number of consensus processes varies based on the unlearning techniques used: a single consensus occurs in parallel unlearning, whereas multiple consensuses are required in sequential unlearning. On the other hand, the computational overheads of different unlearning paradigms on the archive chain are unaffected by model updates.
@@
Here, the transmission cost $C_{tran}$ represents the cost associated with a single operation of uploading $C^{up}_{tran}$ or downloading $C^{down}_{tran}$ a model. The committee consensus cost is $C_{con}$. The cost of a single CH update operation is denoted by $C_{CH}$. Suppose that $L$ blocks and $K$ models need to be updated, including $M$ starting models and $(K-M)$ inherited models.
TABLE II: Comparison between the parallel and sequential unlearning. This includes $L$ blocks and $K$ models that need updating, CH update cost $C_{CH}$, consensus cost $C_{con}$, and transmission cost $C_{tran}$ for uploading or downloading models. These costs are independent of the implementation of the unlearning methods.
Paradigm Computation Cost Update Role
Parallel $(K+L)C_{CH}+$ $C_{con}+2KC_{tran}$ Committee
Sequential $2KC_{CH}+KC_{con}+$ $2KC_{tran}+N^{R}(K-1)C_{tran}$ Client
V-C1 Parallel Unlearning
@@
CH cost. The cost of CH includes the cost of transactions CH $C^{Tx}_{CH}$, and the cost of blocks CH $C^{block}_{CH}$ in the blockchain. Therefore, the CH cost of parallel unlearning is
$C^{Pa}_{CH}=KC^{Tx}_{CH}+NC^{block}_{CH}=(K+L)C_{CH}.$ (19)
@@
Block consensus cost. The committee executes the entire parallel unlearning process, enabling consensus to be reached for multiple transactions or blocks simultaneously. The consensus cost of parallel unlearning is given by
$C^{Pa}_{con}=C_{con}.$ (20)
@@
Transmission cost. The transmission cost of the model is primarily incurred at the consensus nodes for uploading and downloading. Since the consensus nodes can perform uploading and downloading tasks in parallel, the transmission cost of the model is modeled as
$C^{Pa}_{tran}=KC^{up}_{tran}+KC^{down}_{tran}=2KC_{tran}.$ (21)
V-C2 Sequential Unlearning
@@
CH cost. In this process, each transaction update requires one transaction CH and one block CH. The CH cost of sequential unlearning is
$C^{Se}_{CH}=KC^{Tx}_{CH}+KC^{block}_{CH}=2KC_{CH}.$ (22)
@@
Block consensus cost. Each model update requires one round of consensus. The consensus cost for sequential unlearning is
$C^{Se}_{con}=KC_{con}.$ (23)
@@
Transmission cost. The transmission cost of the model primarily involves the uploading and downloading processes by local clients, as well as the downloading process by consensus nodes. Suppose that the average number of inheritances per model is $N^{R}$. The transmission cost for sequential unlearning is given by
$\begin{split}C^{Se}_{tran}&=KC^{up}_{tran}+N^{R}(K-1)C^{down}_{tran}+KC^{down}_{tran}\\ &=2KC_{tran}+N^{R}(K-1)C_{tran}.\end{split}$ (24)
We compare the computational costs and role updates of the two paradigms in Table II. Sequential unlearning incurs higher costs than parallel unlearning. This is because, during model updates, sequential unlearning requires first reaching a consensus and then recording the result on the chain, before proceeding to the next model update. During this process, the costs arise primarily from consensus and transmission overhead. For a model, the computational consumption of its updates can be $2C_{CH}+C_{con}+2C_{tran}+N^{R}C_{tran}$.
In the case of one transaction corresponding to one block and the case of multiple transactions corresponding to one block, the block update costs are denoted by $SC_{block}$ and $MC_{block}$, respectively.
@@
$Tx$-block. One or more transaction updates go through CH overhead $C_{CH}$. Similarly, the main overhead is the same as the transaction overhead for one block or multiple block updates. Suppose that there are $K$ model updates. The costs can be given by
$SC_{block}=K(C^{block}_{CH}+C^{Tx}_{CH})=2KC_{CH}.$ (25)
$Txs$-block. In the case of multiple transactions corresponding to a single block, suppose there are $K$ model updates, then the update cost in this block is
$MC_{block}=C^{block}_{CH}+KC^{Tx}_{CH}=(K+1)C_{CH}.$ (26)
V-C3 Comparison
@@
Based on the comparison of these paradigms, the choice of updating entity significantly affects computational costs. If individual users perform updates, each must upload their updated model to the blockchain and inform the next user, necessitating consensus for each update and incurring transmission overheads. As the number of updated models outpaces the number of block updates, the cumulative cost of CH in the blocks surpasses that of collective updates managed by the committee. Conversely, entrusting the committee with model updates alleviates the computational load on clients. As the number of models increases, consensus and transmission costs are lower than when managed by clients, allowing the committee to focus more on overall system performance and cost efficiency.
VI Experiments
@@
In the experiment, we establish a testbed and assess the effectiveness of the above-mentioned two paradigms–parallel and single unlearning. We evaluate the impact of these paradigms on accuracy for both unlearned and retained labels, incorporating classic models, such as ResNet18 and MobileNetV2. We also evaluate the overhead of interacting with the dual-chain structure. This thorough analysis extends to measuring the time consumption during training and inference, providing insights into the trade-offs between unlearning effectiveness and model utility within the framework.
@@
We align the parallel unlearning paradigm with the gradient ascent and the sequential unlearning paradigm with the re-training. This alignment is reasonable since each method aptly reflects the characteristics of its respective unlearning paradigm: re-training, which requires sequentially fetching and training each model, and gradient ascent, which facilitates parallel processing due to its on-off fetching capability from the blockchain.
VI-A Evaluation Metrics
For illustration convenience, our experiment concentrates on class-level unlearning tasks, although BlockFUL is technically capable of supporting class-, client-, and sample-level unlearning, thanks to the broad compatibility of BlockFUL that supports various prevalent unlearning methods, such as re-training and gradient ascent [56]. We evaluate the model’s accuracy using the training datasets to assess its practicality in experiments. This method is reasonable, as it directly tests whether the removed information continues to influence the model and is a standard practice in the field [57, 58, 44].
@@ VI-A Evaluation Metrics
TABLE III: Unlearning Performance on CIFAR-10
Model #$C_{f}$ Metrics Original Model Re-training Model Gradient Ascent Cumulative Unlearning Time (s)
@@
MobileNetV2 1 $A_{D_{r}}\uparrow$ 81.36% 71.44% 99.40% 79.92% 82.85% 96.83% 51.53% 44.59% 42.02% 181.09 361.77 596.89 28.32 29.9 31.28
$A_{D_{f}}\downarrow$ 77.44% 77.04% 99.98% 0.00% 0.00% 0.00% 7.39% 8.52% 8.60%
2 $A_{D_{r}}\uparrow$ 98.56% 87.71% 82.49% 89.45% 78.76% 87.94% 53.64% 30.56% 17.56% 181.94 376.88 583.02 28.62 29.34 30.09
$A_{D_{f}}\downarrow$ 79.40% 86.21% 78.42% 0.00% 0.00% 0.00% 8.00% 7.16% 5.54%
4 $A_{D_{r}}\uparrow$ 92.43% 86.99% 86.70% 86.15% 94.37% 94.84% 50.97% 35.80% 16.58% 151.14 311.59 471.64 28.66 29.98 32.98
$A_{D_{f}}\downarrow$ 74.46% 74.11% 90.09% 0.00% 0.00% 0.00% 6.55% 9.36% 0.00%
7 $A_{D_{r}}\uparrow$ 92.30% 99.99% 97.39% 99.99% 99.99% 99.29% 51.87% 52.99% 36.53% 84.55 164.17 245.01 28.40 30.38 32.21
$A_{D_{f}}\downarrow$ 90.15% 91.86% 91.68% 0.00% 0.00% 0.00% 8.75% 4.32% 7.90%
ResNet18 1 $A_{D_{r}}\uparrow$ 99.99% 99.99% 89.25% 99.99% 95.85% 99.99% 60.64% 52.78% 54.95% 312.95 625.27 951.14 28.62 30.05 31.59
$A_{D_{f}}\downarrow$ 99.99% 99.99% 92.73% 0.00% 0.00% 0.00% 7.76% 8.38% 6.78%
2 $A_{D_{r}}\uparrow$ 74.63% 90.09% 95.28% 99.99% 94.13% 91.95% 52.22% 60.17% 72.05% 290.76 597.86 907.71 28.67 29.8 31.03
$A_{D_{f}}\downarrow$ 98.82% 90.43% 97.70% 0.00% 0.00% 0.00% 8.95% 5.23% 3.45%
4 $A_{D_{r}}\uparrow$ 80.08% 88.00% 99.87% 90.42% 98.54% 99.95% 41.83% 58.00% 63.92% 218.85 442.17 664.46 28.55 29.86 30.83
$A_{D_{f}}\downarrow$ 89.21% 82.15% 99.68% 0.00% 0.00% 0.00% 4.50% 9.49% 8.50%
7 $A_{D_{r}}\uparrow$ 92.49% 90.19% 98.31% 99.95% 87.81% 99.99% 80.51% 54.49% 48.09% 134.04 264.57 391.35 28.54 30.66 32.27
$A_{D_{f}}\downarrow$ 84.14% 81.98% 97.50% 0.00% 0.00% 0.00% 6.00% 8.52% 7.35%
TABLE IV: Unlearning Performance on Fashion-MNIST
Model #$C_{f}$ Metrics Original Model Re-training Model Gradient Ascent Cumulative Unlearning Time (s)

### v3

@@ 9        Re-training according to (19);
11    Send model_update message($w^{\prime}_{i}$);
[…]
13     $R^{\prime}_{curr}$ $\leftarrow$ Committee.Overwrite(versionversion,
14    newnew versionversion, $R_{curr},CH_{sk}$);
@@ 11    Send model_update message($w^{\prime}_{i}$);
17        if $W^{\prime},v^{\prime}_{i,j},R^{\prime}_{curr}$ isis invalidinvalid then
@@
[…]
@@
Algorithm 2 Sequential unlearning
VI-C Analysis and Comparison
The complexity and overhead analyses of FL and blockchain operations further highlight their computational feasibility and practical implementation.
VI-C1 FL Complexity Analysis
In FL networks, the computational cost of sequential updates increases linearly with the number of requests $Q$, leading to a complexity of $\mathcal{O}\left(QdN^{R}|D|\times|S|\right)$, where $d$ is the depth, $D$ is the dataset size, and $S$ is the number of model parameters. In contrast, parallel updates depend primarily on depth $d$ and remain independent of $Q$, with a complexity of $\mathcal{O}\left(dN^{R}|S|\right)$.
Considering energy consumption, assume there are $L$ blocks and $K$ models to be updated, including $Q$ starting models and (KQ)(K-Q) inherited models. There are $M$ nodes participating in the consensus. The energy cost for sequential updates is $QP_{seq}\times\mathcal{O}\left(dN^{R}|D|\times|S|\right)$, while for parallel updates, it is given by $QP_{seq}\times\mathcal{O}\left(|D|\times|S|\right)+(K-Q)MP_{para}\times\mathcal{O}\left(dN^{R}|S|\right)$, where $P_{seq}$ and $P_{para}$ are the computational power for sequential and parallel updates, respectively.
VI-C2 Blockchain Complexity Analysis
@@
The update is executed every time a node is traversed, regardless of whether or not unlearning has been performed on the node. The number of consensus processes varies based on the unlearning techniques used: a single consensus occurs in parallel unlearning, whereas multiple consensuses are required in sequential unlearning. The computational overheads of the unlearning paradigms on the archive chain are unaffected by model updates.
@@
The transmission cost $C_{tran}$ represents the cost associated with a single operation of uploading $C^{up}_{tran}$ or downloading $C^{down}_{tran}$ a model. The committee consensus cost is $C_{con}$. The cost of a single CH update operation is denoted by $C_{CH}$.
The analysis of parallel unlearning costs is as follows:
@@
CH cost. The cost of CH includes the cost of transactions CH $C^{Tx}_{CH}$, and the cost of blocks CH $C^{block}_{CH}$ in the blockchain. Therefore, the CH cost of parallel unlearning is $C^{Pa}_{CH}=KC^{Tx}_{CH}+NC^{block}_{CH}=(K+L)C_{CH}$.
@@
Block consensus cost. The committee executes the entire parallel unlearning process, enabling consensus to be reached for multiple transactions or blocks simultaneously. The consensus cost of parallel unlearning is $C^{Pa}_{con}=C_{con}$.
@@
Transmission cost. The transmission cost of the model is primarily incurred at the consensus nodes for uploading and downloading. Since the consensus nodes can perform uploading and downloading tasks in parallel, the transmission cost of the model is $C^{Pa}_{tran}=KC^{up}_{tran}+KC^{down}_{tran}=2KC_{tran}$.
For parallel operations, the blockchain overhead is
$(K+L)C_{CH}+C_{con}+2KC_{tran},$ (20)
which can also derive the estimated energy consumption:
$\left(K+L\right)ME_{CH}+ME_{con}+2K|S|E_{tran},$ (21)
where $E_{CH}$ is the CH energy consumption; $E_{con}$ represents the energy consumption per node for consensus among $M$ participants, $E_{tran}$ is the transmission energy consumption per unit of data size, and $S$ is the number of model parameters.
The analysis of sequential unlearning costs is as follows:
@@
CH cost. In this process, each transaction update requires one transaction CH and one block CH. The CH cost of sequential unlearning is $C^{Se}_{CH}=KC^{Tx}_{CH}+KC^{block}_{CH}=2KC_{CH}$.
@@
Block consensus cost. Each model update requires one round of consensus. The consensus cost for sequential unlearning is $C^{Se}_{con}=KC_{con}$.
@@
Transmission cost. The transmission cost of the model primarily involves the uploading and downloading processes by local clients, as well as the downloading process by consensus nodes. Suppose that the average number of inheritances per model is $N^{R}$. The transmission cost for sequential unlearning is $C^{Se}_{tran}=KC^{up}_{tran}+N^{R}(K-1)C^{down}_{tran}+KC^{down}_{tran}=2KC_{tran}+N^{R}(K-1)C_{tran}$.
For sequential operations, the blockchain overhead is
$2KC_{CH}+KC_{con}+2KC_{tran}+N^{R}(K-1)C_{tran},$ (22)
which can also derive the estimated energy consumption:
$2KME_{CH}+KME_{con}+2K|S|E_{tran}+N^{R}\left(K-1\right)|S|E_{tran}.$ (23)
Additionally, we assess other metrics, including blockchain overhead in cases where each transaction forms an individual block ($SC_{block}$) and where multiple transactions are grouped within a single block ($MC_{block}$). We also examine the impact of varying training dataset sizes ($|D|$), considering factors such as depth ($d$) and parameter count ($|S|$).
@@
One or more transaction updates go through CH overhead $C_{CH}$. Similarly, the main overhead is the same as the transaction overhead for one block or multiple block updates. Suppose that there are $K$ model updates. The costs can be given by $SC_{block}=K(C^{block}_{CH}+C^{Tx}_{CH})=2KC_{CH}$. In the case of multiple transactions corresponding to a single block, suppose there are $K$ model updates, then the update cost in this block is $MC_{block}=C^{block}_{CH}+KC^{Tx}_{CH}=(K+1)C_{CH}$.
In large-scale datasets and complex models, computational cost limitations manifest in several ways. First, as dataset size $|D|$ and model complexity (e.g., depth $d$ and parameter count $|S|$) increase, computational costs rise significantly, especially in sequential updates, where complexity grows linearly with the number of requests $Q$, reducing system efficiency. Second, in blockchain systems, the consensus overhead $\mathcal{O}\left(C_{con}\right)$ and on-chain processing overhead $\mathcal{O}\left(C_{CH}\right)$ further add to the computational burden. Finally, while parallel updates can reduce computational complexity, they rely on hardware parallelization, which may not be fully exploited in resource-constrained environments.
VI-C3 Comparison
The design of parallel and sequential unlearning paradigms aligns with real-world applications, where the choice of updating entity significantly impacts computational costs. In financial prediction models, parallel unlearning facilitates rapid updates from multiple market data sources, enabling real-time trend monitoring and dynamic portfolio adjustments to respond effectively to market fluctuations. Conversely, in medical compliance and record management, unlearning is infrequent but critical, often triggered by data lifecycle expiration or regulatory changes, requiring strict compliance and structured execution, making sequential unlearning the preferred approach.
@@ VI-C3 Comparison
The method of executing updates also affects efficiency. If individual users perform updates, each must upload their modified model to the blockchain and notify the next user, necessitating separate consensus processes and incurring significant transmission overhead. As the number of updated models surpasses the number of block updates, the cumulative cost of CH in the blocks exceeds that of collective updates managed by a central committee. Delegating updates to the committee reduces computational overhead on clients, optimizing consensus and transmission costs as model numbers grow. This allows the committee to maintain a more efficient balance between system performance and cost management, ensuring scalable and streamlined unlearning processes.
VII Experiments
@@
In the experiment, we establish a testbed to assess the effectiveness of parallel and sequential unlearning, and evaluate their impact on accuracy for both unlearned and retained labels, incorporating classic models, such as ResNet18 and MobileNetV2. We also evaluate the overhead of interacting with the dual-chain structure.
@@
We align the gradient ascent-based parallel unlearning and re-training-based sequential unlearning. This alignment is reasonable since each method aptly reflects the characteristics of its respective unlearning paradigm: re-training, which requires sequentially fetching and training each model, and gradient ascent, which facilitates parallel processing due to its one-off fetching capability from the live chain [^1] Other unlearning methods, such as fine-tuning and distillation, can also be implemented within these paradigms. For the sake of clarity and focus, we select representative methods for our evaluation in this paper..
VII-A Evaluation Metrics
We focus on class-level unlearning, though BlockFUL supports class-, client-, and sample-level unlearning using methods like re-training and gradient ascent [13, 15]. We evaluate accuracy on training data to check if removed information still affects the model, following standard practice [30].
  • $\bullet$
@@ VII-A Evaluation Metrics
  • $\bullet$
    Blockchain overhead. This includes the time overhead required for CH and consensus in the unlearning process on the blockchain, as well as the time overhead for processing single or multiple transactions within a block.
  • $\bullet$
    Energy consumption. This mainly includes the energy consumption of the FL network model updates and the energy consumption of the blockchain network.
  • $\bullet$
    Latency. This includes the cumulative unlearning time and the time spent on blockchain and CH operations.
TABLE I: Unlearning Performance on CIFAR-10
Model $C_{f}$ Metrics Original (%) Re-training (%) Gradient Ascent (%) Cumulative Unlearning Time (s)
@@
MobileNetV2
1 $A_{D_{r}}\uparrow$ 81.36 71.44 99.40 79.92 82.85 96.83 51.53 44.59 42.02 181.09 361.77 596.89 28.32 29.9 31.28
$A_{D_{f}}\downarrow$ 77.44 77.04 99.98 0.00 0.00 0.00 7.39 8.52 8.60
2 $A_{D_{r}}\uparrow$ 98.56 87.71 82.49 89.45 78.76 87.94 53.64 30.56 17.56 181.94 376.88 583.02 28.62 29.34 30.09
$A_{D_{f}}\downarrow$ 79.40 86.21 78.42 0.00 0.00 0.00 8.00 7.16 5.54
4 $A_{D_{r}}\uparrow$ 92.43 86.99 86.70 86.15 94.37 94.84 50.97 35.80 16.58 151.14 311.59 471.64 28.66 29.98 32.98
$A_{D_{f}}\downarrow$ 74.46 74.11 90.09 0.00 0.00 0.00 6.55 9.36 0.00
ResNet18
1 $A_{D_{r}}\uparrow$ 99.99 99.99 89.25 99.99 95.85 99.99 60.64 52.78 54.95 312.95 625.27 951.14 28.62 30.05 31.59
$A_{D_{f}}\downarrow$ 99.99 99.99 92.73 0.00 0.00 0.00 7.76 8.38 6.78
2 $A_{D_{r}}\uparrow$ 74.63 90.09 95.28 99.99 94.13 91.95 52.22 60.17 72.05 290.76 597.86 907.71 28.67 29.8 31.03
$A_{D_{f}}\downarrow$ 98.82 90.43 97.70 0.00 0.00 0.00 8.95 5.23 3.45
4 $A_{D_{r}}\uparrow$ 80.08 88.00 99.87 90.42 98.54 99.95 41.83 58.00 63.92 218.85 442.17 664.46 28.55 29.86 30.83
$A_{D_{f}}\downarrow$ 89.21 82.15 99.68 0.00 0.00 0.00 4.50 9.49 8.50
TABLE II: Unlearning Performance on Fashion-MNIST
Model $C_{f}$ Metrics Original (%) Re-training (%) Gradient Ascent (%) Cumulative Unlearning Time (s)